A 63kg client has been told to cycle at 60% of her 7 MET capacity. What absolute VO2 and Workrate (W) will she be exercising at?
Answer #1

One metabolic equivalent of task (MET) = 3.5 ml oxygen per kg per minute.

VO2 = 7 × 3.5 => 24.5

Work rate = 3.5 × MET × Body weight (in Kg)

=> 3.5 × 7 × 63 = answer 1543.5 watts.

Calories burned/min = 1543.5/200 => 7.71 calories per minute.

hope it helps.